# DP-2Stage: Adapting Language Models as Differentially Private Tabular Data Generators

### v1

@@ 5.3 DP Benchmarks
Table 3 presents a comparison of DP-2Stage-O against GAN and VAE-based methods under arguably strict privacy setting of $\varepsilon=\{1,8\}$. Each model was trained five times, with results averaged across runs and four synthetic datasets generated per run. The evaluation was conducted using tabular data metrics described in section 5.1. Our proposed methods, DP-2Stage-U and DP-2Stage-O, demonstrate competitive performance across all metrics on both datasets. Notably, DP-CTGAN and DP-GAN show greater improvement on the Adult dataset at $\varepsilon=1$, while DP-2Stage-O outperforms DP-GAN at $\varepsilon=8$. However, scaling DP-GAN to higher $\varepsilon$ values for the Airline dataset proved challenging, requiring up to five days to run before the process was terminated.
Adult Airline
Method F1 AUC ACC HIST F1 AUC ACC HIST
$\varepsilon=1$ DP-GAN 33.5$\pm$20 67.7$\pm$9 64.2$\pm$10 63.7$\pm$3 40.2$\pm$24 63.9$\pm$13 59.8$\pm$6 44.7$\pm$12
@@
$\varepsilon=8$ DP-GAN 19.6$\pm$20 50.0$\pm$0 50.0$\pm$26 33.3$\pm$9 - - - -
@@
Table 3: DP Benchmark. Utility metrics (F1, AUC, and ACC) are presented as the averages of logistic regression and XGBoost performance. HIST represents the average histogram intersection scores calculated using bins of 20 and 50. Results are averaged across five model runs and four synthetic datasets per run with standard deviation reported after $\pm$. The best value per column for each $\varepsilon$ is shown in bold while second best value is underlined.
@@ 5.4 DP-2Stage Benchmark
In Table 3, we present the results of our proposed DP-2Stage method, which utilizes two pseudo-dataset approaches (DP-2Stage-U and DP-2Stage-O) as described in Section 4.1. The results demonstrate consistent improvements in downstream task utility, with DP-2Stage-O achieving the best accuracy metrics for the Adult dataset and DP-2Stage-U outperforming for the Airline dataset. Furthermore, DP-2Stage-U, leveraging uniform pseudo-data, achieves the fastest inference time—up to 21x faster (see Table 7 in Appendix F)—making it particularly advantageous for real-world deployment.
@@ 5.4.1 Impact of Column Shuffling
Next, we evaluate the impact of column shuffling. In Figure 3, we compared the perplexity of DP-2Stage method with DP-Standard. The top section shows the evolution of perplexity with column shuffling enabled, the default in prior work Borisov et al. (2023), while in the bottom we shows perplexity without column shuffling. The results demonstrate that DP-2Stage consistently outperforms DP-Standard in perplexity across both settings, with the non-shuffling configuration yielding the best performance under DP.
In Table 4, we provide a detailed comparison of these two approaches across the evaluated datasets. The results suggest that column shuffling generally benefits the Non-DP setting, which may explains its frequent use in prior research. However, under DP, models without shuffling more often achieve better performance, as indicated by the underlined or bold values, although histogram metrics occasionally favor models with shuffling. Additionally, enabling shuffling significantly increases sampling time (approximately 42-51 hours per model run), particularly for DP-Standard and DP-2Stage-O model (see Table 7 in Appendix F), as a result, we only present model trained on a single run for DP-Standard and DP-2Stage-O to ensure a fair evaluation when shuffling is enabled.
@@
DP-Standard 3.537 24.6$\pm$23 62.4$\pm$8 63.9$\pm$14 87.4$\pm$0 4.276 44.8$\pm$33 73.9$\pm$12 66.9$\pm$11 93.3$\pm$1
3.224 31.4$\pm$17 62.1$\pm$8 68.4$\pm$5 85.9$\pm$1 3.790 59.4$\pm$6 65.8$\pm$4 61.8$\pm$2 88.3$\pm$1
$\varepsilon=1$ DP-2Stage-U 2.973$\pm$0.01 19.7$\pm$13 48.0$\pm$8 61.9$\pm$13 87.6$\pm$1 3.932$\pm$0.08 48.8$\pm$11 57.5$\pm$7 55.2$\pm$5 93.4$\pm$1
@@
DP-2Stage-O 3.270 22.7$\pm$22 55.1$\pm$10 68.0$\pm$9 87.1$\pm$1 3.948 59.8$\pm$4 68.9$\pm$6 66.3$\pm$5 92.0$\pm$1
3.049 27.8$\pm$18 60.2$\pm$7 65.5$\pm$9 88.4$\pm$0 3.737 73.2$\pm$2 82.0$\pm$2 75.7$\pm$2 91.7$\pm$1
Table 4: Comparison of Methods with Shuffle Enabled () or Disabled (). The best result under DP is highlighted in bold, while the top-performing result for each shuffle setting is underlined. Non-DP methods generally perform comparably across both settings but tend to show better results with shuffle enabled (). Conversely, DP methods often achieve higher performance when shuffle is disabled (). Results are averaged across five model runs, each using a different random seed, with four synthetic datasets generated per run. For DP-Standard and DP-2Stage-O, results are based on a single model run, averaging across four synthetic datasets, due to the high inference time required with shuffle enabled (approximately 42–51 hours per run; see Table 7 in Appendix F), ensuring a fair evaluation.
@@ 5.4.2 Impact of Weighted Loss
To evaluate the impact of weighting the loss in the second stage (see Equation 6), we conducted experiments using both the default loss ($\lambda=0.33$) and the proposed weighting value ($\lambda=0.65$). The results, presented in Table 5, indicate that emphasizing value tokens during the second-stage DP fine-tuning yields the best performance. This approach also improves the performance of DP-Standard, though the improvement is more pronounced for DP-2Stage-O.
Despite the gains observed in DP-Standard when $\lambda=0.65$, DP-2Stage offers the better performance across both settings. On the Adult dataset, DP-2Stage-U performs slightly better with the default loss in most cases (except for the F1 score), whereas on the Airline dataset, it performs worse with the default loss.
@@
DP-2Stage-O default 30.3$\pm$15 60.4$\pm$7 65.8$\pm$9 87.3$\pm$1 48.8$\pm$17 55.6$\pm$19 54.4$\pm$15 92.4$\pm$1
$0.65$ 30.4$\pm$17 61.6$\pm$8 66.7$\pm$8 88.5$\pm$1 55.2$\pm$18 62.5$\pm$19 60.0$\pm$16 92.5$\pm$1
@@
5.4.3 Impact of Low Perplexity on Utility metrics
In Table 4, we present perplexity values with and without column shuffling alongside utility performance metrics. While uniform pseudo-data (DP-2Stage-U) often achieves lower perplexity, this does not always translate to improved downstream utility or histogram scores. In some cases, the out-distribution pseudo-data (DP-2Stage-O) approach yields better performance. We hypothesize that this discrepancy observed in DP-2Stage-U may result from overfitting when pseudo-data closely resembles the private data. In such cases, the model may lack sufficient incentive to learn meaningful correlations in the private data during the second stage. Conversely, when out-distribution pseudo-data is used, the model is encouraged to learn these correlations while retaining the structural information established during stage 1. This approach may provide better overall performance compared to directly applying DP.
@@ 6 Limitations & Future Work
Incorporating Additional Datasets.
The results presented in our work are based on two datasets, Adult and Airline, which differ in domain, column size, and sample size. Expanding the analysis to include a wider variety of datasets would allow for a more thorough evaluation and enhance the validity of the claims made in this paper. However, due to the computational cost associated with training and inference, we leave this for future work.
Metric Diversity.
Our evaluation primarily used perplexity to assess the language model’s performance, machine learning efficacy to measure the downstream task utility, and Histogram Intersection to evaluate the synthetic data fidelity. While these metrics are informative and widely adopted in synthetic tabular data research (Xu et al., 2019; Afonja et al., 2023), they do not provide a complete picture of synthetic data quality. Incorporating domain-specific metrics could offer deeper insights into the data’s applicability and reveal limitations not captured by general metrics (Chen et al., 2024).
@@ Acknowledgement
This work is supported by ELSA – European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617, Bundesministeriums fur Bildung und Forschung (PriSyn), grant No. 16KISAO29K, and Medizininformatik-Plattform “Privatsphärenschutzende Analytik in der Medizin” (PrivateAIM), grant No. 01ZZ2316G. Computation resources used in this work are supported by the Helmholtz Association’s Initiative and Networking Fund on the HAICORE@FZJ partition and CISPA Helmholtz Center for Information Security computing services. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them. We thank Joscha Cüppers for providing detailed feedback on this work.
@@ Appendix A Differential Privacy: more details
A.1 Privacy Analysis
@@ C.1 Perplexity
  • Low Perplexity: Suggests that the model is confident in its predictions and is performing well, predicting the next word accurately.
@@ D.1 Datasets
    Adult Income Dataset columns: Age, Work Class, FNLWGT, Education, Education Number, Marital Status, Occupation, Relationship, Race, Sex, Capital Gain, Capital Loss, Hours per Week, Native Country, and Income
@@
    Airline Passenger Satisfaction Dataset columns: ID, Gender, Customer Type, Age, Type of Travel, Class, Flight Distance, Inflight Wi-Fi Service, Departure/Arrival Time Convenience, Ease of Online Booking, Gate Location, Food and Drink, Online Boarding, Seat Comfort, Inflight Entertainment, Onboard Service, Leg Room Service, Baggage Handling, Check-in Service, Inflight Service, Cleanliness, Departure Delay (minutes), Arrival Delay (minutes), and Satisfaction (Neutral or Dissatisfied, Satisfied).
@@
Appendix E Extended Results
Table 6 presents the results for the machine learning models evaluated: XGBoost (XGB) and Logistic Regression (LR). Histogram Intersection score (HIST) is reported for two bin sizes: 20 and 50. The averaged values are summarized in Table 3.
Dataset
$\varepsilon=1$ Adult DP-Standard 13.9$\pm$7 55.5$\pm$6 73.0$\pm$1 41.6$\pm$5 61.4$\pm$7 57.4$\pm$6 85.1$\pm$2 86.2$\pm$2
DP-2Stage-O 15.0$\pm$7 55.9$\pm$6 73.5$\pm$1 45.9$\pm$5 67.3$\pm$6 59.8$\pm$5 88.2$\pm$1 88.8$\pm$1
Airline DP-Standard 55.8$\pm$3 62.1$\pm$7 60.0$\pm$6 65.1$\pm$6 68.4$\pm$10 64.9$\pm$8 89.6$\pm$4 91.1$\pm$3
DP-2Stage-O 53.5$\pm$16 61.8$\pm$16 59.2$\pm$13 56.9$\pm$20 63.3$\pm$22 60.8$\pm$18 92.0$\pm$1 93.0$\pm$1
$\varepsilon=8$ Adult DP-Standard 17.3$\pm$6 58.6$\pm$4 73.6$\pm$1 45.2$\pm$6 65.8$\pm$8 61.8$\pm$4 84.1$\pm$1 84.9$\pm$1
DP-2Stage-U 9.7$\pm$6 49.4$\pm$4 74.4$\pm$1 35.1$\pm$8 54.1$\pm$9 52.9$\pm$5 86.4$\pm$1 87.4$\pm$1
Table 6: DP-GPT-2 Benchmarks. For each dataset, the best value corresponding to different privacy budgets ($\varepsilon$) is highlighted in bold. Results are averaged across five model runs with varying random seeds, with four synthetic datasets generated per run. LR refers to the Logistic Regression model, and XGB represents the XGBoost model.
Appendix F Ablation
F.1 Sampling time

### v2

@@ 5.3 DP Benchmarks
Table 3 presents a comparison of DP-2Stage against GAN and VAE-based methods under strict privacy setting of $\varepsilon=1$. Our proposed methods, DP-2Stage-U and DP-2Stage-O, demonstrate competitive performance across all metrics on all datasets. Notably, DP-CTGAN and DP-GAN show greater utility metric improvements on the Adult dataset but underperform on fidelity metrics (CorAcc, Pair, and Hist). These metrics are crucial for assessing the model’s ability to preserve statistical relationships in real data, including marginal distributions and pairwise correlations.
In Appendix F.1, Table 7, we present results for a higher privacy budget ($\varepsilon=8$). As the privacy budget increases – meaning privacy constraints are relaxed – we observe corresponding improvements in the utility metrics for DP-2Stage.
Dataset Method F1 AUC ACC CorAcc Pair HIST
Adult Non-LLM
$\varepsilon=1,$ DP-GAN 33.5$\pm$20 67.7$\pm$9 64.2$\pm$10 39.9$\pm$3 41.2$\pm$4 63.7$\pm$3
$\delta=10^{-5}$ DP-CTGAN 42.2$\pm$20 78.0$\pm$7 75.7$\pm$3 51.3$\pm$3 59.2$\pm$2 75.7$\pm$2
DP-VAE 0.0$\pm$0 50.0$\pm$0 75.6$\pm$0 48.8$\pm$1 40.3$\pm$1 61.8$\pm$2
@@
DP-Standard 27.8$\pm$15 58.5$\pm$7 65.2$\pm$9 55.0$\pm$1 68.4$\pm$1 85.7$\pm$2
DP-2Stage-U 21.2$\pm$12 48.9$\pm$6 61.9$\pm$13 55.0$\pm$1 76.1$\pm$1 86.7$\pm$1
DP-2Stage-O
+airline 30.4$\pm$17 61.6$\pm$8 66.7$\pm$8 55.4$\pm$1 72.3$\pm$1 88.5$\pm$1
+texas 31.6$\pm$13 60.5$\pm$7 66.4$\pm$8 55.6$\pm$1 71.3$\pm$1 86.9$\pm$1
Airline Non-LLM
$\varepsilon=1,$ DP-GAN 40.2$\pm$24 63.9$\pm$13 59.8$\pm$6 37.4$\pm$9 22.2$\pm$13 44.7$\pm$12
$\delta=10^{-5}$ DP-CTGAN 67.1$\pm$8 76.7$\pm$8 68.0$\pm$6 31.7$\pm$2 62.2$\pm$2 78.7$\pm$2
DP-VAE 26.5$\pm$28 57.9$\pm$13 57.3$\pm$6 46.6$\pm$1 20.6$\pm$0 41.8$\pm$1
@@
DP-Standard 60.5$\pm$7 65.3$\pm$9 62.4$\pm$7 64.0$\pm$2 77.0$\pm$2 90.3$\pm$3
DP-2Stage-U 68.5$\pm$9 77.8$\pm$10 72.1$\pm$7 65.3$\pm$1 80.8$\pm$1 90.7$\pm$1
DP-2Stage-O
+adult 55.2$\pm$18 62.5$\pm$19 60.0$\pm$16 66.8$\pm$1 80.1$\pm$1 92.5$\pm$1
+texas 52.5$\pm$13 61.0$\pm$13 58.4$\pm$10 66.1$\pm$2 78.7$\pm$2 90.4$\pm$1
Texas Non-LLM
$\varepsilon=1,$ DP-GAN 13.7$\pm$16 58.3$\pm$12 78.3$\pm$8 36.1$\pm$7 34.6$\pm$5 68.3$\pm$7
$\delta=10^{-5}$ DP-CTGAN 63.9$\pm$11 91.4$\pm$3 82.6$\pm$19 43.9$\pm$6 66.9$\pm$6 84.7$\pm$3
DP-VAE 0.0$\pm$0 50.0$\pm$0 82.5$\pm$0 62.1$\pm$1 43.9$\pm$1 77.9$\pm$1
GPT-2
DP-Standard 55.4$\pm$10 90.2$\pm$5 77.1$\pm$11 70.3$\pm$1 60.6$\pm$1 92.3$\pm$1
DP-2Stage-U 23.5$\pm$14 59.8$\pm$14 67.3$\pm$17 68.2$\pm$0 80.7$\pm$6 93.4$\pm$0
DP-2Stage-O
+adult 74.8$\pm$4 96.7$\pm$1 89.1$\pm$3 69.9$\pm$2 60.5$\pm$2 91.7$\pm$1
+airline 74.3$\pm$5 96.4$\pm$0 88.8$\pm$3 70.9$\pm$1 62.0$\pm$2 93.2$\pm$0
Table 3: DP Benchmark. Utility metrics (F1, AUC, and ACC) are presented as the averages of two ML Models (XGBoost and Logistic Regression). Pair, and Hist are reported as averages of two bin sizes (Bins 20 and 50). Results are averaged across five model runs and four synthetic datasets per run with standard deviation reported after $\pm$. The best value per column for $\varepsilon=1$ is shown in bold while second best value is underlined.
@@ 5.4 DP-2Stage Benchmark
We compare our proposed DP-2Stage methods, which leverage two pseudo-dataset approaches (DP-2Stage-U and DP-2Stage-O) as described in Section 4.1, against DP-Standard, which directly fine-tunes the LLM under a strict privacy budget ($\varepsilon=1$). Our methods consistently improve downstream task utility and fidelity. As shown in Table 3, DP-2Stage-O achieves the highest F1, AUC, and ACC scores on the Adult and Texas datasets but performs slightly worse on the Airline dataset compared to DP-2Stage-U and DP-Standard. We hypothesize that the choice of pseudo data may have influenced this outcome – Adult and Texas datasets may not provide sufficient structural information for effective second-stage DP fine-tuning on Airline. Further investigation is needed to understand the key properties required in pseudo data to optimize performance. Additionally, our approaches achieve the best results across all fidelity metrics i.e CorAcc, Pair, and Hist.
@@ 5.4.1 Impact of Column Shuffling
Next, we evaluate the impact of column shuffling. In Figure 3, we compare the perplexity of DP and Non-DP LLM models on the Adult dataset. The top section illustrates the evolution of perplexity with column shuffling enabled – the default setting in prior work by Borisov et al. (2023) – while the bottom section shows perplexity without shuffling. The results indicate that the non-shuffling configuration yields the lowest perplexity under DP.
Table 4 presents a detailed comparison of these two approaches on the Adult and Airline datasets. The findings suggest that column shuffling generally benefits the Non-DP setting, which may explain its widespread use in prior research. However, under DP, models without shuffling tend to achieve better performance, as highlighted by the bold values. This advantage may stem from the added complexity that shuffling introduces to DP training, potentially amplifying the effects of gradient perturbation [^11] When column shuffling is applied during DP training, it adds complexity to the learning process. This additional complexity could make it harder for the model to learn meaningful patterns, especially under DP constraints, where noise is deliberately added to gradients to ensure privacy. Because DP already introduces randomness into the training process, shuffling may further disrupt the model’s ability to learn effectively by increasing variability in the data order. This, in turn, could amplify the impact of the noise added by DP, making training less stable and reducing performance.. That said, the Hist metric show better performance with shuffling except for DP-2Stage-O on Adult dataset.
Additionally, enabling shuffling significantly increases sampling time (approximately 42–51 hours per model run), particularly for DP-Standard and DP-2Stage-O (see Table 6 in Appendix E). As a result, to ensure a fair evaluation when shuffling is enabled, we report results from a single training run for DP-Standard and DP-2Stage-O.
@@
GPT-2 DP-Standard 3.537 24.6$\pm$23 62.4$\pm$8 63.9$\pm$14 87.4$\pm$0 4.276 44.8$\pm$33 73.9$\pm$12 66.9$\pm$11 93.3$\pm$1
$\varepsilon=1,$ 3.224 31.4$\pm$17 62.1$\pm$8 68.4$\pm$5 85.9$\pm$1 3.790 59.4$\pm$6 65.8$\pm$4 61.8$\pm$2 88.3$\pm$1
$\delta=10^{-5}$
DP-2Stage-U 2.973$\pm$0.01 19.7$\pm$13 48.0$\pm$8 61.9$\pm$13 87.6$\pm$1 3.932$\pm$0.08 48.8$\pm$11 57.5$\pm$7 55.2$\pm$5 93.4$\pm$1
@@
DP-2Stage-O
+adult - - - - - 3.948 59.8$\pm$4 68.9$\pm$6 66.3$\pm$5 92.0$\pm$1
- - - - - 3.737 73.2$\pm$2 82.0$\pm$2 75.7$\pm$2 91.7$\pm$1
+airline 3.270 22.7$\pm$22 55.1$\pm$10 68.0$\pm$9 87.1$\pm$1 - - - - -
3.049 27.8$\pm$18 60.2$\pm$7 65.5$\pm$9 88.4$\pm$0 - - - - -
Table 4: Comparison of Methods with Shuffle Enabled () or Disabled (). The best result under DP is highlighted in bold, while the top-performing result for each shuffle setting is underlined. Non-DP methods generally perform comparably across both settings but tend to show better results with shuffle enabled (). Conversely, DP methods often achieve higher performance when shuffle is disabled (). Results are averaged across five model runs, each using a different random seed, with four synthetic datasets generated per run. For DP-Standard and DP-2Stage-O, results are based on a single model run, averaging across four synthetic datasets, due to the high inference time required with shuffle enabled (approximately 42–51 hours per run; see Appendix E, Table 6), ensuring a fair evaluation.
@@ 5.4.2 Impact of Weighted Loss
To assess the impact of loss weighting in the second stage (see Equation 6), we conducted experiments using both the default setting ($\lambda=0.33$) and the proposed weighting value ($\lambda=0.65$). As shown in Table 5, emphasizing value tokens during second-stage DP fine-tuning leads to the best overall performance. This approach also improves DP-Standard, though the gains are more pronounced for DP-2Stage-O.
On the Adult dataset, DP-2Stage-U generally performs better with the default loss, except for the F1 score, whereas on the Airline dataset, it performs worse with the default loss. These results suggest that loss weighting benefits both DP-2Stage-O and DP-Standard, though its effects on DP-2Stage-U remain inconsistent, warranting further investigation.
@@
DP-2Stage-O
+adult default - - - - 48.8$\pm$17 55.6$\pm$19 54.4$\pm$15 92.4$\pm$1
$0.65$ 55.2$\pm$18 62.5$\pm$19 60.0$\pm$16 92.5$\pm$1
+airline default 30.3$\pm$15 60.4$\pm$7 65.8$\pm$9 87.3$\pm$1 - - - -
$0.65$ 30.4$\pm$17 61.6$\pm$8 66.7$\pm$8 88.5$\pm$1 - - - -
@@
5.4.3 Understanding the Utility-Perplexity Gap in DP-2Stage-U
In Table 4, we present perplexity values with and without column shuffling alongside utility performance metrics. While uniform pseudo-data (DP-2Stage-U) often achieves lower perplexity, this does not always translate to improved downstream utility or histogram scores. In some cases, the out-distribution pseudo-data (DP-2Stage-O) approach yields better performance. We hypothesize that this discrepancy observed in DP-2Stage-U may result from overfitting when pseudo-data closely resembles the private data. In such cases, the model may lack sufficient incentive to learn meaningful correlations in the private data during the second stage. Conversely, when out-distribution pseudo-data is used (DP-2Stage-O), the model is encouraged to learn these correlations while retaining the structural information established during stage 1. This approach may provide better overall performance compared to directly applying DP. Notably, DP-2Stage-U achieves the fastest inference time – up to 21x faster (see Appendix E, Table 6) – making it particularly advantageous for real-world deployment.
5.4.4 Marginal-based DP-Baselines.
In Appendix F.2, we present results for two marginal-based DP methods: MST (McKenna et al., 2021) and AIM (McKenna et al., 2022). As shown in Appendix F.2, Table 8, these methods achieve strong performance across most metrics. However, they underperform on certain fidelity metrics, such as Pair and Hist, particularly for the Airline dataset – likely due to the large number of numerical columns in this dataset.
@@ 6 Limitations & Future Work
Privacy Guarantees (in Stage 1).
In Stage 1, assuming that the public dataset used is indeed “public”, the privacy guarantees of our method hold exactly, as only Stage 2 is exposed to private data and is trained with differential privacy. This is the case for our DP-2Stage-O method.
However, in Stage 1, the proposed DP-2Stage-U method utilizes private data to compute statistics for the uniformly sampled dataset, which weakens the stated privacy guarantees. To ensure that the privacy guarantees from the second stage hold completely, these statistics should instead be computed in a differentially private manner by allocating a small privacy budget. Future work should explore this approach.
@@ Acknowledgement
This work is supported by ELSA – European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617, Bundesministeriums fur Bildung und Forschung (PriSyn), grant No. 16KISAO29K, and Medizininformatik-Plattform “Privatsphärenschutzende Analytik in der Medizin” (PrivateAIM), grant No. 01ZZ2316G. Computation resources used in this work are supported by the Helmholtz Association’s Initiative and Networking Fund on the HAICORE@FZJ partition and CISPA Helmholtz Center for Information Security computing services. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them. We thank the TMLR reviewers and action editor for their detailed feedback, which has undoubtedly improved the quality of our work. We also thank Joscha Cüppers for providing valuable feedback on this work.
@@ Appendix A Differential Privacy: more details
A.1 Standard DP
We define the Standard DP or DP-Standard training process as fine-tuning the LLM directly on private data using the Differentially Private Stochastic Gradient Descent (DPSGD) (Abadi et al., 2016) mechanism. DPSGD ensures that the training process complies with the formal definition of differential privacy (refer to Definition 3.2) through the following steps: (1) Gradients are computed for each individual sample within a mini-batch. (2) Gradients are clipped to a fixed norm to bound sensitivity. (3) Gaussian noise, calibrated to the privacy parameters ($\varepsilon,\delta$), is added to the aggregated gradients. (4) The resulting noisy gradients are used to update the model’s parameters.
A.2 Privacy Analysis
@@ C.1 Perplexity
  • Low Perplexity: Suggests that the model is confident in its predictions and is performing well, predicting the next token accurately.
@@ C.2.2 Normalized Histogram Intersection
C.2.3 Pairwise Correlation Similarity Accuracy (CorAcc)
We evaluate the correlation between data columns using the approach described by  Tao et al. (2021) and  Afonja et al. (2023). Specifically, we use Cramer’s V with bias correction for categorical columns, the Correlation Ratio for numerical-categorical columns, and the Pearson Correlation Coefficient (absolute values) for numerical columns. The ranges for these measures are as follows: Cramer’s V and Correlation Ratio are bounded between 0 and 1, while the Pearson Correlation Coefficient spans -1 to 1. Following  Tao et al. (2021), correlation values are discretized into four levels: low [0, 0.1), weak [0.1, 0.3), medium [0.3, 0.5), and strong [0.5, 1). The CorAcc metric quantifies the similarity between synthetic and original data by measuring the fraction of column pairs where the assigned correlation levels match.
C.2.4 Pairwise Attribute Distribution Similarity (Pair)
This metric extends the Normalized Histogram Intersection (HIST) by calculating the histogram intersection for all two-way marginals and averaging the results across all attribute pairs. For numerical columns, we discretize the values into bins of size 20 and 50 before computing the intersections.
@@ D.1 Datasets
Texas Dataset.
The Texas Hospital Discharge dataset [^12] https://www.dshs.texas.gov/thcic/ is a large public use data file provided by the Texas Department of State Health Services. We used the preprocessed version which consists of 100,000 records uniformly selected from a pre-processed file containing patient data from 2013131313https://github.com/spring-epfl/synthetic_data_release/blob/master/data/texas.csv version from Stadler et al. (2022). We retain 18 attributes and assume a binary classification task by predicting only minor and major mortality risk following the setup of  Afonja et al. (2023). Duplicates where also removed. The final size of the dataset was therefore reduced to 75,105 which was split to non-overlapping train/test/validation. Validation size is fixed to 1000 for all dataset.
No additional preprocessing was done for Adult and Airline other than removing duplicates.
List of Column names:
@@
    Adult Income : Age, Work Class, FNLWGT, Education, Education Number, Marital Status, Occupation, Relationship, Race, Sex, Capital Gain, Capital Loss, Hours per Week, Native Country, and Income
@@
    Airline Passenger Satisfaction: ID, Gender, Customer Type, Age, Type of Travel, Class, Flight Distance, Inflight Wi-Fi Service, Departure/Arrival Time Convenience, Ease of Online Booking, Gate Location, Food and Drink, Online Boarding, Seat Comfort, Inflight Entertainment, Onboard Service, Leg Room Service, Baggage Handling, Check-in Service, Inflight Service, Cleanliness, Departure Delay (minutes), Arrival Delay (minutes), and Satisfaction (Neutral or Dissatisfied, Satisfied).
  3. 3.
    Texas: Discharge, Type of Admission, Patient State, Patient Status, Sex Code, Race, Ethnicity, Admission Weekday, Patient Age, Illness Severity, Length of Stay, Total Charges, Total Non-Covered Charges, Total Charges for Accommodation, Total Non-Covered Charges for Accommodation, Total Charges for Ancillary Services, Total Non-Covered Charges for Ancillary Services, and Risk of Mortality.
@@
Appendix E Sampling Time